OpenStudy (xxlyh):

How many solutions can be found for the equation x + 5 = x? Zero One Two Infinitely many

OpenStudy (xxlyh):

Colton solved an equation incorrectly, as shown below: Step 1: x – 13 = 26 Step 2: x = 26 – 13 Step 3: x = 39 Which statement best explains why Step 2 is incorrect in Colton's solution? He did not subtract 13 from 26. He did not multiply 26 by 13. He did not add 13 to 26. He did not divide 26 by 13.

OpenStudy (xxlyh):

last one xDD sorrryy omg

OpenStudy (threshsupport):

zero

OpenStudy (anonymous):

for the first one the only solution I could think of is that you just add 5 to x.

OpenStudy (xxlyh):

@AihberKhan I got these wrong in a test and they didn't show the correct answer so I was wonderingg

OpenStudy (anonymous):

so I think it must be 1

OpenStudy (anonymous):

so you are just reviewing these or something like that?

OpenStudy (skullpatrol):

x + 5 = x is saying take any number and add 5 to it do you get the same number?

OpenStudy (anonymous):

see x + 5 = ? Plug in any number for it

OpenStudy (xxlyh):

yup im just reviewing^^

OpenStudy (anonymous):

btw look at the second step in the first question

OpenStudy (anonymous):

do you see anything wrong with the subtraction

OpenStudy (anonymous):

and how much she says it equals

OpenStudy (anonymous):

or he

OpenStudy (xxlyh):

ohhh yeahh

OpenStudy (anonymous):

and how many solutions to x + 5 can you think of

OpenStudy (xxlyh):

I'm kinda confused with that one

OpenStudy (anonymous):

imagine 1 + 5 = ? How many ways are there to add 1 + 5

OpenStudy (xxlyh):

1?

OpenStudy (anonymous):

there are probably millions of ways or maybe billions but there is only 1 way unless you swap it around

OpenStudy (skullpatrol):

What number, call it x, will make x + 5 = x A true statement?

OpenStudy (anonymous):

there is one way

OpenStudy (anonymous):

and that is through thinking

OpenStudy (anonymous):

I forgot that it was x + 5 = x sorry.

OpenStudy (xxlyh):

its okay xD

OpenStudy (xxlyh):

what about the second question?

OpenStudy (anonymous):

if you have 5 + ? = 5 what is the other number

OpenStudy (anonymous):

for the other number remember when I said they forgot to add and they instead subtracted

OpenStudy (anonymous):

for number 2

OpenStudy (anonymous):

instead of 26 + 13 they did 26 - 13

OpenStudy (anonymous):

also if you wanted to check just go up a few replies

OpenStudy (xxlyh):

oh okie i get it now

OpenStudy (anonymous):

but for the first one

OpenStudy (anonymous):

imagine 5 + ? = 5

OpenStudy (anonymous):

what is the other number

OpenStudy (xxlyh):

0?

OpenStudy (anonymous):

yup

OpenStudy (xxlyh):

ohh so thats the answer for the first one?

OpenStudy (skullpatrol):

Let x = any real number Then x + 5 = 5 more than x x =/= x + 5 Only x = x.

OpenStudy (anonymous):

yea 0 and they forgot to subtracted and didn't add for the second one

OpenStudy (anonymous):

they forgot to add and subtracted

OpenStudy (xxlyh):

thank you for all your help!!

OpenStudy (anonymous):

so instead of 13 + 26 they did 26 - 13

OpenStudy (anonymous):

np